Relationships between Kp,Kc and Kx for reactions​ involving ideal gases
1
Expert's answer
2017-05-04T06:13:29-0400
The relationships between Kp, Kc and Kx for reaction: aA + bB = cC + dD are:
KP = KC (RT) ^(c+d–a–b ) = KX P^( c+d–a–b)